One diagonal of a parallelogram is 17 cm long and makes angles of 36° and 15° with the sides. Find the lengths (x and y) of the sides.
Answer by stanbon(75887)

One diagonal of a parallelogram is 17 cm long and makes angles of 36° and 15° with the sides. Find the lengths (x and y) of the sides.
--------------
The 3rd angle is 180-(36+15) = 129 degrees
---
Law of sines:
x/sin(36) = 17/sin(129)
x = 12.86
------
Law of Sines:
y/sin(15) = 17/sin(129)
y = 5.66
-----
Cheers,
Stan H.
========